### v1

# AAAI Press Formatting Instructions for Authors Using — A Guide

Abstract
Reconstructing objects and extracting high-quality surfaces play a vital role in the real world. Current 4D representations show the ability to render high-quality novel views for dynamic objects but cannot reconstruct high-quality meshes due to their implicit or geometrically inaccurate representations. In this paper, we propose a novel representation that can reconstruct accurate meshes from sparse image input, named Dynamic 2D Gaussians (D-2DGS). We adopt 2D Gaussians for basic geometry representation and use sparse-controlled points to capture 2D Gaussian’s deformation. By extracting the object mask from the rendered high-quality image and masking the rendered depth map, a high-quality dynamic mesh sequence of the object can be extracted. Experiments demonstrate that our D-2DGS is outstanding in reconstructing high-quality meshes from sparse input. More demos and code are available at https://github.com/hustvl/Dynamic-2DGS.
00footnotetext: *Equal contributions.   🖂Corresponding author.
Introduction
The natural world is dynamic, and how to extract the geometric shape of dynamic objects from the sparse input images is an important research problem in computer vision. The main challenges lie in the complex and non-rigid motion, which makes it hard to model accurate motion. Some 3D representations  (Wang et al. 2023b; Huang et al. 2024a) succeed in reconstructing high-quality mesh from static objects but may fall short in 4D objects since we cannot leave every life stock still.
Many approaches succeed in representing dynamic scenes and rendering high-quality novel views. Dynamic NeRFs (Pumarola et al. 2021; Park et al. 2021a, b) mainly suffer from their implicit representations which cause unfriendly memory consumption. Dynamic Gaussian splatting (Wu et al. 2024; Yang et al. 2023a; Huang et al. 2024b) representations can maintain a set of explicit 3D Gaussians (Kerbl et al. 2023) for efficient NVS with high training efficiency, the proposed deformation field shows the potential to modeling object’s motion accurately. However, 3D Gaussians are mainly for NVS and do not have multiview consistency, which causes inaccurate geometry (Huang et al. 2024a). We believe that a proper representation that can reconstruct dynamic objects may achieve two goals: (a) modeling complex motion with high efficiency and (b) enabling the export of high-quality and smooth geometry.
Recently, 2D Gaussian Splatting (2DGS) (Huang et al. 2024a) uses a collection of two-dimensional oriented planar Gaussian disks to represent scenes, ensuring consistent view geometry when modeling surfaces. Extending 2D Gaussians to dynamic object reconstruction and get a high-quality surface is a natural but difficult topic. There are two challenges and the first problem is How to extract a clean consistent surface from sparse input? Geometry floaters also exist in the deformation of 2D Gaussian primitives: the Gaussian owns the same color as the background which is hard to be pruned and would induce inaccurate depth, triggering the low-quality mesh extracted by Truncated Signed Distance Function (TSDF). Moreover, modeling accurate object’s motion also remains a hard issue in the dynamic representations. It’s hard to preserve correspondence between 2D Gaussians’s motion and real-world deformation, which may lead to unstable surfaces. We also observe that while 2D Gaussians benefit from multiview consistency due to their isotropic characteristic, it also degrades Gaussian’s fitting ability when dealing with sparse views.
To tackle the aforementioned problems, we propose to use the sparse-controlled points method (Huang et al. 2024b) to model 2D Gaussians’ motion. Our findings indicate that sparse-controlled points accurately model semi-rigid motion because nearby 2D Gaussian surfels are determined by the sparse-controlled points which provide a more precise canonical-world transformation, ensuring the geometric accuracy of 2D Gaussians in the canonical space and result in smoother surfaces, as illustrated in Fig. LABEL:fig:teaserfig. Furthermore, we introduce a filtering method to remove geometry floaters by filtering the depth image using the rendered RGB mask. This approach allows for an accurate surface mesh to be obtained via the TSDF method.
Our contributions can be summarized as follows:
  • We propose Dynamic 2D Gaussians (D-2DGS), a novel framework that employs sparse-controlled points to guide 2D Gaussians’ deformation and reconstruct the accurate dynamic mesh. Our framework takes into account both the dynamic and geometric properties of objects.
  • To remove Geometry floaters, we propose to filter the depth image using a mask extracted from the rendered high-quality RGB image.
  • Experiments demonstrate that our D-2DGS achieves state-of-the-art (SOTA) reconstruction quality compared with other advanced representations.
Figure 1: Illustration of Geometry floaters. In the canonical space, all the spatial Gaussian primitives are connected by the control points. After deformation, both objects Gaussians and floater Gaussians are transformed by control points. The similar color to the background makes it hard to be pruned, which causes inaccurate geometry.
Related Works
3D/4D Representations
Representing the 4D world and rendering photo-realistic novel view images has been an important and challenging topic. Explicit representations such as mesh (Collet et al. 2015; Li et al. 2018; Guo et al. 2015; Su et al. 2020), voxels (Guo et al. 2019; Hu et al. 2022; Li et al. 2017), and point clouds enjoy editable manipulations and GPU-friendly applications but suffer from strict multi-view constraints or more sensors. Recently, implicit (Martin-Brualla et al. 2021; Park et al. 2021a, b) representations have demonstrated their efficiency for novel view synthesis from sparse input and many approaches (Xu et al. 2022; Fridovich-Keil et al. 2023; Cao and Johnson 2023; Wang et al. 2023a; Shao et al. 2023; Gan et al. 2023; Li et al. 2022) combine the benefits of these representations. Notably, Gaussian Splatting (Kerbl et al. 2023; Yang et al. 2023b; Wu et al. 2024; Yang et al. 2023a; Zhang et al. 2024) (GS) based representations as an efficient mixture presentation, show the ability to render high-quality novel views and support many downstream tasks (Jiang et al. 2024; Ji et al. 2024; Yi et al. 2024a, b; Yin et al. 2023; Liang et al. 2024). Recent 4D representation, SC-GS (Huang et al. 2024b), proposes a sparse control points method to control the deformation of 3D Gaussians, which reaches outstanding NVS quality and high training efficiency. However, most 3D/4D representations are only for NVS or extract mesh from static scenes and there is less research focus on extracting accurate mesh from dynamic objects in monocular input. Our D-2DGS shows the ability to reconstruct the accurate 3D mesh from 2D images captured in the 4D real world with high efficiency.
Geometry Reconstruction
Geometry reconstructions are mainly from streamable multiview (Collet et al. 2015; Li et al. 2018; Guo et al. 2015; Su et al. 2020) constraints. Recently, 3DGS (Kerbl et al. 2023) is proven to enjoy high training efficiency and shows great potential to reconstruct high-quality mesh from spatial Gaussians. Many works are proposed to improve the geometric accuracy of 3DGS. SuGaR (Guédon and Lepetit 2024) introduces a regularization term that encourages the Gaussians to align well with the surface of the scene. NeuSG (Chen, Li, and Lee 2023) jointly optimizes 3DGS and neural implicit models, combining the advantages of both, enabling the generation of complete surfaces with high details. Similarly, 3DGSR (Lyu et al. 2024) also encourages the alignment of 3D Gaussian and SDF to improve the surface reconstruction effect. 2D Gaussian Splatting (2DGS) (Huang et al. 2024a)is a breakthrough in mesh reconstruction since the view geometry consistency is maintained by planar Gaussian disks.
The above geometric reconstruction methods are all for static scenes. Recently, some works have been attempted on the geometric reconstruction of dynamic scenes. Dynamic Gaussian Mesh (DG-Mesh) (Liu, Su, and Wang 2024) aims to reconstruct a high-fidelity and temporally consistent mesh from a given monocular video. DG-Mesh introduces Gaussian grid anchoring to encourage uniformly distributed Gaussian distributions and obtains better grid reconstruction by grid densification and pruning of deformed Gaussian distributions. MaGS (Ma, Luo, and Yang 2024) propose mutual-adsorbed mesh Gaussian representation to refine the generated mesh. Vidu4D (Wang et al. 2024b) uses Dynamic Gaussian Surfels as scene representation primitives, achieving text to 4D generation with high-quality appearance and geometry. Our methods propose to model Gaussian surfels’ motion by sparse-controlled points, maintaining consistent and smooth surface mesh generation.
Preliminaries
2D Gaussian Splatting
2D Gaussian Splatting (2DGS) simplifies 3D modeling through 2D Gaussians. Each 2D Gaussians is represented by a central point $P_{k}$, two principal tangential vectors $\mathbf{t}_{u}$ and $\mathbf{t}_{v}$, and a scaling vector $\mathbf{S}=(s_{u},s_{v})$ that controls the variance of the 2D Gaussian. The original normal of the 2D Gaussian is defined by two orthogonal tangent vectors $\mathbf{t}_{w}=\mathbf{t}_{u}\times\mathbf{t}_{v}$. The 2D Gaussian function can be parameterized on the local tangent plane in world space as:
$P(u,v)=P_{k}+s_{u}\mathbf{t}_{u}u+s_{v}\mathbf{t}_{v}v.$ (1)
For a point $\mathbf{u}=(u,v)$ in the $uv$ space, its 2D Gaussian value is:
$\mathcal{G}(\mathbf{u})=exp(-\frac{u^{2}+v^{2}}{2}).$ (2)
The parameters of each 2D Gaussian also include opacity $\alpha$ and view-dependent appearance $c$ parameterized with spherical harmonics.
2D Gaussians are sorted by their center depth and organized into tiles based on their bounding boxes. The alpha-weighted appearance is integrated from front to back using volumetric alpha blending:
$c(x)=\sum_{i=1}\mathbf{c}_{i}\alpha\mathcal{G}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-\alpha_{j}\mathcal{G}(\mathbf{u}(x))).$ (3)
However, 2D Gaussians can only reconstruct 3D mesh from static scenes. We propose to use Sparse-Controlled points to model 2D Gaussians deformation and extend 2D Gaussians representing the 4D world successfully.
Sparse-Controlled Gaussian Splatting
Sparse-Controlled Gaussian Splatting (SC-GS) (Huang et al. 2024b) shows excellent performance in the task of synthesizing novel views in dynamic scenes. It uses sparse control points to control dense 3D Gaussians in the scene. For each control point at timestep $t$, SC-GS uses MLP $\Phi$ to predict a rotation matrix $R_{i}^{t}\in\mathbf{SE}(3)$ and rotation matrix $T_{i}^{t}\in\mathbb{R}^{3}$. For each Gaussian $G_{j}:(\mu_{j},q_{j},s_{j},\sigma_{j},sh_{j})$, use k-nearest neighbor (KNN) search to get the $K$ neighboring control points $\{p_{i},i=0,1,...,k\}$. SC-GS employs LBS (Sumner, Schmid, and Pauly 2007) $\mathcal{I}$ to compute the warped Gaussian $\mu_{j}^{t}$ and $q_{j}^{t}$:
$\mu_{j}^{t}=\sum_{k\in\mathcal{N}_{j}}w_{jk}(R_{k}^{t}(\mu_{j}-p_{k}))+p_{k}+T_{k}^{t}),$ (4)
$q_{j}^{t}=(\sum_{k\in\mathcal{N}_{j}}w_{jk}r_{k}^{t})\otimes q_{j},$ (5)
where $R_{k}^{t}$ and $r_{k}^{t}$ are the matrix and quaternion representations of the predicted rotation at control point $k$, respectively. The weight $w_{jk}$ is calculated by:
$w_{jk}=\frac{\hat{w}_{jk}}{\sum_{k\in\mathcal{N}_{j}}\hat{w}_{jk}},\mathrm{where}~{}\hat{w}_{jk}=exp(-\frac{d_{jk}^{2}}{2o_{k}^{2}}),$ (6)
where $d_{jk}$ is the distance between the Gaussian and the adjacent control point, and $o_{k}$ is the learned radius parameter of the control point.
However, SC-GS (Huang et al. 2024b) mainly chooses 3D-GS (Kerbl et al. 2023) as its geometry representation, leading to multiview inconsistency. Meanwhile, There are fewer geometry constraints during the optimization of canonical Gaussians, we mainly choose 2DGS (Huang et al. 2024a) as basic geometry, and propose an outlier removal method to overcome the Geometry floaters as shown in Fig. 1 from the sparse monocular input.
Method
Figure 2: Framework of our D-2DGS. Sparse points are bonded with canonical 2D Gaussians. Deformation networks are used to predict each sparse control point’s control signals given any timestamp. The image and depth are rendered by deformed 2D Gaussians with alpha blending. To get high-quality meshes, depth images are filtered by rendered images with RGB mask, and then TSDF is applied on multiview depth images and RGB images.
Dynamic 2D Gaussians Framework
The pipeline of our framework is illustrated in Fig. 2. We adopt 2D Gaussian primitives $\mathcal{G}$ for scene representation. We follow deformation-based Gaussian splatting to compute the deformation of each 2D Gaussians at timestamp $t_{j}$:
$\mathcal{G}^{j}=Deform(\mathcal{G},t_{j},\mathbf{P};\Phi),$ (7)
where $Deform$ denotes deformation process, $\mathbf{P}=\{p_{i}\}_{i=1}^{N}$ denotes sparse control points. Then $\mathcal{G}^{j}$ are used to render the RGB image $\mathbf{I}$ and the depth image $\mathbf{D}$ by differential splatting algorithm (Huang et al. 2024b; Yifan et al. 2019) $\mathcal{S}$ with a camera parameters $[\mathbf{R},\mathbf{t}]$:
$\mathbf{I},\mathbf{D}=\mathcal{S}\left(\mathcal{G}^{j};\theta),[\mathbf{R},\mathbf{t}]\right).$ (8)
We then extract a mask $\mathbf{M}$ from the rendered image $\mathbf{I}$ and apply it to the depth image $\mathbf{D}$ to filter out floaters:
$\mathbf{D}^{\prime}=Flitering(\mathbf{I},\mathbf{D}),$ (9)
the $Flitering$ progress will be discussed later.
Finally, the refined depth image $\mathbf{D}^{\prime}$ is used to extract the mesh using the Truncated Signed Distance Function (TSDF).
Sparse point controlled 2D Gaussians
To accurately extract dynamic mesh sequences, it is essential to account for both the dynamic and geometric characteristics of the modeled objects. We leverage the multi-view geometric consistency of 2D Gaussian representations, selecting 2D Gaussians $\mathcal{G}:(\mu,q,s,\sigma,\textit{sh})$ as the foundational primitive for scene representation.
The accuracy of motion is ensured by sparse control points which are proposed in SC-GS (Huang et al. 2024b). The sparse controlled points $\mathbf{P}$ and deformation network $\Phi(p_{i};\theta)$ are proposed to model the accurate deformation. Where $p_{i}$ represents the position of the $i$-th control point and $\theta$ denotes the parameters of the deformation network. The control signals at $t_{j}$ are computed by the MLP $\Phi$:
$[R_{i}^{j},T_{i}^{j}]=\Phi(p_{i},t_{j}).$ (10)
Given $K$ adjacent control points $\mathbf{p}=\{p_{0},\dots,p_{k}\}$, the 6-DoF transform $\{R^{j},T^{j}\}$ can be computed by a deformation network $\Phi$ applying on $p$, and the deformed 2D Gaussians $\mathcal{G}^{j}$ at $t_{j}$ can be interpolated using the LBS (Sumner, Schmid, and Pauly 2007) $\mathcal{I}$ including Eq. 4 and Eq. 5 :
$\mathcal{G}^{j}=\mathcal{I}\left(\{R^{j},T^{j}\},\mathcal{G}\right).$ (11)
This approach ensures that both the temporal dynamics and geometric consistency are preserved, facilitating the extraction of high-fidelity dynamic mesh sequences.
Figure 3: Mesh and mesh rendering visualization results of the D-NeRF dataset.
Method Lego Bouncingballs Jumpingjacks Hook
[…]
DG-mesh* 23.474 0.903 0.103 28.045 0.972 0.088 31.667 0.977 0.036 29.217 0.965 0.049
D-3DGS 22.534 0.884 0.115 23.667 0.956 0.102 28.878 0.972 0.039 27.499 0.959 0.048
SCGS 22.784 0.882 0.113 25.231 0.956 0.103 29.534 0.976 0.035 27.095 0.961 0.048
D-2DGS (Ours) 23.293 0.887 0.112 27.786 0.969 0.073 29.293 0.974 0.032 27.802 0.962 0.042
Method Mutant Standup Trex Hellwarrior
[…]
DG-mesh* 31.369 0.970 0.043 32.220 0.979 0.036 29.559 0.970 0.044 27.602 0.971 0.055
D-3DGS 28.342 0.957 0.048 29.946 0.977 0.031 28.061 0.958 0.054 25.986 0.965 0.053
SCGS 28.002 0.955 0.047 29.273 0.977 0.032 28.521 0.962 0.050 26.419 0.967 0.048
D-2DGS (Ours) 28.120 0.960 0.042 29.512 0.976 0.028 28.677 0.967 0.043 25.441 0.958 0.051
Table 1: Rendering quality of extracted meshes on the D-NeRF (Pumarola et al. 2021) dataset. The color of each cell means: best, second best. (*DG-mesh needs to query additional MLPs when rendering mesh.)
Depth filtering
During the optimization process, Geometry floaters manifest as numerous background-colored artifacts, despite the high quality of the rendered image. In the depth image, this issue is evident as noise points located outside the actual object boundaries. Consequently, after mesh extraction, a significant number of floating artifacts appear around the object.
To mitigate these artifacts, we propose a masking technique $Flitering$ as mentioned in Eq. 9 that involves extracting the object’s mask from the rendered high-quality image and applying it to filter the depth image. The process can be formalized as follows:
The extraction of the mask $\mathbf{M}$ from the rendered image $\mathbf{I}$ can be mathematically described as follows:
$\mathbf{M}(x)=\mathbf{1}[\mathbf{I}(x)=bg],$ (12)
where $\mathbf{1}[\cdot]$ is the indicator function that returns 1 if the color of $\mathbf{I}(x)$ is different from the background color $bg$, and 0 otherwise.
The mask $\mathbf{M}$ is then used to refine the depth image $\mathbf{D}$ by element-wise multiplication:
$\mathbf{D}^{\prime}=\mathbf{D}\odot\mathbf{M},$ (13)
where $\odot$ represents the element-wise product. The resulting filtered depth image $\mathbf{D}^{\prime}$ effectively removes the floating noise points, leading to a cleaner mesh extraction with reduced artifacts.
Loss Function
In order to normalize the spatial distribution of Gaussians, we introduce the Depth Distortion and Normal Consistency regularization terms in 2DGS. The Depth Distortion regularization term encourages Gaussians to be distributed at the same depth as much as possible, as shown in the following formula:
$L_{d}=\sum_{i,j}\omega_{i}\omega_{j}\left|z_{i}-z_{j}\right|,$ (14)
where $i$ indexes over intersected splats along the ray, $\omega_{i}$ denotes the blending weight of the intersection point. $\omega_{i}=\alpha_{i}\mathcal{G}_{i}(u(x)){\textstyle\prod_{j=1}^{i-1}}(1-\alpha\mathcal{G}_{j}(u(x)))$ $z_{i}$ is the depth of the intersection points.
The Normal Consistency regularization term can make the normals of 2D Gaussians as close as possible to the normals of the object:
$L_{n}=\sum_{i}\omega_{i}(1-n_{i}^{\mathrm{T}}N),$ (15)
where $n_{i}$ represents the normal of Gaussians, $N$ normal estimated by the nearby depth point $p$, it is calculated by the following formula:
$N(x,y)=\frac{\nabla_{x}p\times\nabla_{y}p}{\left|\nabla_{x}p\times\nabla_{y}p\right|}.$ (16)
Therefore, the overall loss function is:
$L_{all}=L_{1}(I_{r},I_{g})+\lambda_{s}L_{ssim}(I_{r},I_{g})+\lambda_{n}L_{n}+\lambda_{d}L_{d},$ (17)
where $I_{r}$ represents the rendered image, and $I_{g}$ represents the Ground Truth. $\lambda_{s}$ and $\lambda_{n}$ are the weights of each loss function respectively.
Method D-NeRF datasets DG-Mesh datasets
[…]
DG-Mesh* 29.114 0.963 0.057 0.698 0.123
D-3DGS 26.864 0.954 0.061 0.895 0.115
SCGS 27.107 0.954 0.060 0.856 0.115
D-2DGS(ours) 27.490 0.957 0.052 0.851 0.132
Table 2: Average results on both D-NeRF datasets and DG-Mesh datasets. (*DG-mesh needs to query additional MLPs when rendering mesh.)
Mesh Extraction
The mesh at any normalized timestamp $t_{i}\in[0,1]$ can be extracted from the trained model. Given $t_{i}$, the model renders multi-view images $\mathbf{I}(t_{i})$ and depth images $\mathbf{D}(t_{i})$. The mesh is then obtained by applying TSDF using Open3D (Zhou, Park, and Koltun 2018):
$\text{Mesh}=\text{TSDF}\left(\{\mathbf{I}_{i},\mathbf{D}_{i}^{\prime}\mid i\in T\}\right).$ (18)
Experiment
Experimental Setups
We mainly compare some novel view synthesis methods for dynamic scenes based on Gaussian Splatting, including DG-Mesh (Liu, Su, and Wang 2024), Deformable 3DGS (Yang et al. 2023a), and SCGS (Huang et al. 2024b). We evaluate the quality of meshes extracted by these methods on the D-NeRF (Pumarola et al. 2021) and DG-Mesh (Liu, Su, and Wang 2024) datasets. Both datasets have large and non-rigid motions and belong to efficient multiview setups as referred to the Dycheck (Gao et al. 2022). Each timestamp has only one view which is randomly selected. Among them, the D-NeRF dataset does not contain the ground truth of the mesh but the DG-Mesh dataset includes the GT mesh. Therefore, We report the PSNR, LPIPS, and SSIM to evaluate the rendered images in the D-NeRF datasets and use Chamfer Distance (CD) and Earth Mover’s Distance (EMD) as mesh evaluation metrics. The mesh extracted by deformable 3DGS and SC-GS following the same settings as ours and 2D-GS (Huang et al. 2024a), and it is worth noting that when DG-Mesh renders an image, two additional MLPs should be maintained when querying the color of the mesh.
Figure 4: Mesh and mesh rendering visualization results of the DG-Mesh dataset.
Implementation Details
We build dynamic 2DGS upon the open-source SCGS (Huang et al. 2024b) codebase. Training iterations for all scenarios are set to 80,000. The training time will take 1 to 2 hours. The number of control points is set to 1024. All experiments are conducted on an NVIDIA RTX A5000 GPU. The hyperparameters $\lambda_{s}$, $\lambda_{n}$ and $\lambda_{d}$ are 1, 0.02 and 1000 respectively. The other hyperparameters are the same as SCGS. When the mesh is extracted, the original dynamic 2DGS representations can be removed.
Method GS render image Mesh
Color Smooth Detail
DG-Mesh
D-3DGS
SCGS
Ours
Table 3: The comparison of different models.
[…]
w/o Depth filtering 0.425 0.110
w/o Normal Consistency 0.443 0.101
w/o Depth Distortion 0.350 0.109
Full model 0.328 0.110
Table 4: Ablation experiment results of the “bird” sample in the DG-Mesh dataset.
Results
The results on the D-NeRF (Pumarola et al. 2021) dataset are shown in Fig. 3. The ideal surface should be as smooth as possible while still containing the details of the object. The mesh extracted by DG-Mesh (Liu, Su, and Wang 2024) has a low resolution and lacks some details of objects, such as the hands of the ‘jumpingjacks’. Deformable-3DGS and SCGS both use 3D Gaussians as the basic representational primitives, which lack geometric accuracy. The extracted mesh overfits the details and is not geometrically accurate and smooth, such as the heads of the ‘jumpingjacks’. The three comparative methods are unable to simulate smooth planar and spherical surfaces, as demonstrated by the “bouncingballs” example. The sphere appears broken and the chassis is uneven due to the influence of lighting. In comparison, our Dynamic 2DGS highlights reconstructing rich details and modeling smooth surfaces.
The rendering results of the mesh are shown in Tab. 1. When DG-Mesh renders an image, it needs to query additional MLPs to obtain the color of the mesh, thus achieving higher PSNR and SSIM. The proposed D-2DGS achieved higher indicators compared with D-3DGS and SCGS.
As shown in Fig. 4, on the DG-Mesh dataset, the proposed method can extract smoother meshes compared to D-3DGS and SCGS. At the same time, compared to DG-Mesh, the proposed method can show the details of objects, such as the mouth of the ‘horse’. The average metrics of the D-NeRF and DG-Mesh datasets are shown in Tab. 2. The data in the DG-Mesh dataset has less details, so the proposed model is comparable to other models in terms of metrics.
We summarize the methods on the Tab. 3. DG-Mesh produces low-quality rendered images through Gaussian rendering, and the extracted meshes rely on MLP modeling for color, and lacking in detail. Meanwhile, D-3DGS and SCGS extract meshes that are over-fitted to details, resulting in surfaces that are not sufficiently smooth.
Figure 5: Ablation Study on Sparse point control.
Ablation Study
Sparse point control.
We conducted ablation experiments on sparse point control, where we removed the sparse point control and directly modeled the offset for 2DGS using an MLP as the offset field. As shown in Fig. 5, in areas with significant motion, such as a person’s hand and a horse’s hoof, sparse point control allows for the extraction of a more detailed mesh.
Figure 6: Ablation Study on Depth filtering.
Depth filtering.
We removed the mask on the Depth image and then conducted experiments. As can be seen in Fig. 6 (a), there are many noise points around the object in the rendered depth image, which may be caused by the offset field offsetting the 2D Gaussians. After extracting the mesh through TSDF fusion, some floating objects will appear around the mesh of the object. With the help of high-quality renderings, the object mask is extracted and the depth image is masked. The noise in the depth image can be filtered out, thereby removing the floating objects in the mesh, as shown in Fig. 6 (b).
Loss function.
We conduct ablation experiments on various loss functions to verify the effectiveness of the loss functions. As shown in Fig. 7, the two regularization terms in the loss function are important for extracting high-quality meshes. The Normal Consistency regularization term can make the normals of Gaussians more accurate, and the Depth Distortion term can make the Gaussians distributed on the plane. The corresponding metrics are shown in Tab. 4.
Limitations and Future work
Though dynamic 2D Gaussians can model the accurate dynamic surface, there are several limitations: (1) Some holes exist in the extracted mesh. (2) Given the extremely sparse views, such as strict monocular input, robustness priors can other low-rank motion representations (Wang et al. 2024a) could be explored to model more accurate motion. (3) More post-processing methods could be extended to handle Geometry floaters and holes completion.
Figure 7: Ablation Study on loss function.
Conclusion
In this paper, we propose Dynamic 2D Gaussians to extract high-quality meshes from dynamic 2D images input. We use 2D Gaussians as the basic representation primitives and use sparse control points to control 2D Gaussians to model the motion of objects. We used high-quality rendered images to filter out the noise in the depth image and used TSDF to obtain higher-quality meshes. Experiments show that the mesh extracted by the proposed method has high details and a smoother surface.
Appendix
Written by AAAI Press Staff1With help from the AAAI Publications Committee.
AAAI Style Contributions by Pater Patel Schneider, Sunil Issar,
[…]
Figure 8: Relighting and composition results in generated mesh on the Blender.
Figure 9: Rendered images, normals, and depth maps.
Figure 10: Comparison of results with different background colors.
In the supplementary material, we mainly provide additional experimental results in Sec. 1. Then more discussions are conducted in Sec. 2. Finally, we delve into the limitations of our proposed Dynamic 2D Gaussians in Sec. 3.
Additional Experimental Results
In the D-NeRF dataset, the metrics of the RGB images rendered from the extracted meshes do not fully reflect the quality of the meshes. Therefore, we provide additional visualization results, as shown in Fig. 11. DG-Mesh exhibits relatively poorer performance in modeling object details. The meshes extracted by Deformable-3DGS and SCGS exhibit numerous pits and holes on the surface. In comparison, the meshes extracted by the proposed Dynamic 2D Gaussians approach capture object details while maintaining a smoother surface, reflecting better geometric properties.
Figure 11: Mesh and mesh rendering visualization results of the D-NeRF dataset.
In the DG-Mesh dataset, the detailed metrics are shown in Table 1. Since the data of the DGMesh dataset is relatively simple and the surface is smooth, it is difficult to reflect the advantages of the proposed method in terms of metrics. Our metrics are close to those of SCGS, but the visual quality shows that our extracted surfaces appear smoother than SCGS. The DG-Mesh method shows advantages in the Duck and Torus2sphere examples, but performs relatively poorly on the other cases. Although DGMesh has a high metrics in Torus2sphere, from the visualization results in Fig. 12, DGMesh does not show a particularly good reconstruction effect in the Torus2sphere example.
The proposed method can render Gaussians to obtain normal depth maps and high-quality RGB images of dynamic objects, as shown in Fig. 9.
Method Duck Horse Bird
[…]
D-3DGS 1.064 0.076 0.298 0.114 0.427 0.123
SCGS 1.001 0.074 0.232 0.118 0.336 0.121
DG-mesh 0.782 0.048 0.299 0.168 0.557 0.128
Ours 1.040 0.092 0.391 0.177 0.328 0.110
Method Beagle Torus2sphere Girlwalk
[…]
D-3DGS 0.585 0.101 2.655 0.155 0.340 0.122
SCGS 0.528 0.100 2.716 0.155 0.321 0.124
DG-mesh 0.626 0.114 1.515 0.126 0.406 0.153
Ours 0.544 0.122 2.479 0.164 0.324 0.129
Table 5: Mesh reconstruction quality comparison results of DG-Mesh dataset.
Figure 12: Visualization results of DGMesh dataset.
Discussion and Application
Background Color.
Training models with different background colors can have some influence on the extraction of the mesh. The background color should be selected to have significant contrast with the color of the object itself. For the training images, the background color should be selected to have significant contrast with the color of the object itself. As shown in Fig. 10, if a black background is used, the shadows on the object are difficult to distinguish from the background, making it challenging to extract the object’s mask from the rendered images. This then makes it difficult to filter out floating artifacts in the depth maps, ultimately leading to the presence of many floating artifacts in the extracted meshes.
Mesh composition and relighting.
The mesh reconstructed by Dynamic 2D Gaussians can be applied in the Blender in the Fig. 8. Placing a proper light source, the shadow exists in the image due to the occlusion relationship between the objects.
Time-Aware Mesh.
Our method supports the extraction of dynamic mesh sequences, as shown in Fig. 13. The extracted precise temporal geometry provides the potential for analyzing the motion characteristics and dynamics of objects.
Figure 13: Time-Aware Mesh.
Limitations
Due to the thin nature of 2DGS and the sparsity of dynamic scene views, the extracted mesh may be broken in some areas. As shown in Fig. 14, the mesh is prone to damage in the following two situations: 1. In some places that are difficult to observe from the training perspective and in shadows, the mesh is prone to damage, such as Fig. 14 (a). 2. Mesh is prone to damage on uneven thin planes with uneven lighting, such as Fig. 14 (b). Subsequent research is needed on better methods for representing precise dynamic geometry, as well as targeted mesh damage repair techniques. This will help us better extract mesh from dynamic objects.
Figure 14: Some areas of the mesh will appear broken.

### v2

@@
-AAAI Press Formatting Instructions for Authors Using — A Guide
-Abstract
+Dynamic 2D Gaussians: Geometrically Accurate Radiance Fields for Dynamic Objects
+Abstract.
 Reconstructing objects and extracting high-quality surfaces play a vital role in the real world.
-Current 4D representations show the ability to render high-quality novel views for dynamic objects but cannot reconstruct high-quality meshes due to their implicit or geometrically inaccurate representations.
+Current 4D representations show the ability to render high-quality novel views for dynamic objects, but cannot reconstruct high-quality meshes due to their implicit or geometrically inaccurate representations.
 In this paper, we propose a novel representation that can reconstruct accurate meshes from sparse image input, named Dynamic 2D Gaussians (D-2DGS).
-We adopt 2D Gaussians for basic geometry representation and use sparse-controlled points to capture 2D Gaussian’s deformation.
-By extracting the object mask from the rendered high-quality image and masking the rendered depth map, a high-quality dynamic mesh sequence of the object can be extracted.
-Experiments demonstrate that our D-2DGS is outstanding in reconstructing high-quality meshes from sparse input.
-More demos and code are available at https://github.com/hustvl/Dynamic-2DGS.
-00footnotetext: *Equal contributions.   🖂Corresponding author.
-Introduction
+We adopt 2D Gaussians for basic geometry representation and use sparse-controlled points to capture the 2D Gaussian’s deformation.
+By extracting the object mask from the rendered high-quality image and masking the rendered depth map, we remove floaters that are prone to occur during reconstruction and can extract high-quality dynamic mesh sequences of dynamic objects.
+Experiments demonstrate that our D-2DGS is outstanding in reconstructing detailed and smooth high-quality meshes from sparse inputs.
+The code is available at https://github.com/hustvl/Dynamic-2DGS.
+Gaussian Splatting; 4D Reconstruction; Sparse Perspective
+submissionid: 3241journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755328isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer graphics
+Figure 1. Applications of D-2DGS. Given sparse 2D dynamic images, D-2DGS can reconstruct dynamic objects and extract high-quality normal, time-aware mesh, render quality depth, and novel view images.
+1. Introduction
 The natural world is dynamic, and how to extract the geometric shape of dynamic objects from the sparse input images is an important research problem in computer vision.
-The main challenges lie in the complex and non-rigid motion, which makes it hard to model accurate motion.
-Some 3D representations  (Wang et al.
-2023b; Huang et al.
-2024a) succeed in reconstructing high-quality mesh from static objects but may fall short in 4D objects since we cannot leave every life stock still.
+The main challenges lie in the complex and non-rigid motion, which makes it hard to model motion accurately.
+Some 3D representations  (Wang et al., 2023; Huang et al., 2024b) succeed in reconstructing high-quality mesh from static objects but may fall short in 4D objects since we cannot leave every life stock still.
 Many approaches succeed in representing dynamic scenes and rendering high-quality novel views.
-Dynamic NeRFs (Pumarola et al.
-2021; Park et al.
-2021a, b) mainly suffer from their implicit representations which cause unfriendly memory consumption.
-Dynamic Gaussian splatting (Wu et al.
-2024; Yang et al.
-2023a; Huang et al.
-2024b) representations can maintain a set of explicit 3D Gaussians (Kerbl et al.
-2023) for efficient NVS with high training efficiency, the proposed deformation field shows the potential to modeling object’s motion accurately.
-However, 3D Gaussians are mainly for NVS and do not have multiview consistency, which causes inaccurate geometry (Huang et al.
-2024a).
+Dynamic NeRFs (Pumarola et al., 2021; Park et al., 2021a, b) mainly suffer from their implicit representations, which cause unfriendly memory consumption.
+Dynamic Gaussian splatting (Wu et al., 2024; Yang et al., 2024a; Huang et al., 2024a) representations can maintain a set of explicit 3D Gaussians (Kerbl et al., 2023) for efficient NVS with high training efficiency, and the proposed deformation field shows the potential to model an object’s motion accurately.
+However, 3D Gaussians are mainly for NVS and do not have multiview consistency, which causes inaccurate geometry (Huang et al., 2024b).
+Some 3D Gaussian-based approaches also succeed in reconstructing a 3D mesh from dynamic images.
+We also found it (Liu et al., 2025) lacks high-quality details.
 We believe that a proper representation that can reconstruct dynamic objects may achieve two goals: (a) modeling complex motion with high efficiency and (b) enabling the export of high-quality and smooth geometry.
-Recently, 2D Gaussian Splatting (2DGS) (Huang et al.
-2024a) uses a collection of two-dimensional oriented planar Gaussian disks to represent scenes, ensuring consistent view geometry when modeling surfaces.
+Recently, 2D Gaussian Splatting (2DGS) (Huang et al., 2024b) uses a collection of two-dimensional oriented planar Gaussian disks to represent scenes, ensuring consistent view geometry when modeling surfaces.
 Extending 2D Gaussians to dynamic object reconstruction and get a high-quality surface is a natural but difficult topic.
-There are two challenges and the first problem is How to extract a clean consistent surface from sparse input?
-Geometry floaters also exist in the deformation of 2D Gaussian primitives: the Gaussian owns the same color as the background which is hard to be pruned and would induce inaccurate depth, triggering the low-quality mesh extracted by Truncated Signed Distance Function (TSDF).
-Moreover, modeling accurate object’s motion also remains a hard issue in the dynamic representations.
-It’s hard to preserve correspondence between 2D Gaussians’s motion and real-world deformation, which may lead to unstable surfaces.
-We also observe that while 2D Gaussians benefit from multiview consistency due to their isotropic characteristic, it also degrades Gaussian’s fitting ability when dealing with sparse views.
-To tackle the aforementioned problems, we propose to use the sparse-controlled points method (Huang et al.
-2024b) to model 2D Gaussians’ motion.
-Our findings indicate that sparse-controlled points accurately model semi-rigid motion because nearby 2D Gaussian surfels are determined by the sparse-controlled points which provide a more precise canonical-world transformation, ensuring the geometric accuracy of 2D Gaussians in the canonical space and result in smoother surfaces, as illustrated in Fig.
-LABEL:fig:teaserfig.
-Furthermore, we introduce a filtering method to remove geometry floaters by filtering the depth image using the rendered RGB mask.
+There are two challenges, and the first problem is How to extract a clean, consistent surface from sparse input?
+Geometry floaters also exist in the deformation of 2D Gaussian primitives: the Gaussian owns the same color as the background, which is hard to be pruned and would induce inaccurate depth, triggering the low-quality mesh extracted by Truncated Signed Distance Function (TSDF).
+Moreover, modeling an accurate object’s motion also remains a hard issue in the dynamic representations.
+It’s hard to preserve correspondence between 2D Gaussians’ motion and real-world deformation, which may lead to unstable surfaces.
+Meanwhile, We observe that while 2D Gaussians benefit from multiview consistency due to their isotropic characteristic, it also degrades the Gaussian’s fitting ability when dealing with sparse views.
+To tackle the aforementioned problems, we propose to use the sparse-controlled points method (Huang et al., 2024a) to model 2D Gaussians’ motion.
+Our findings indicate that sparse-controlled points accurately model semi-rigid motion because nearby 2D Gaussian surfels are determined by the sparse-controlled points, which provide a more precise canonical-world transformation, ensuring the geometric accuracy of 2D Gaussians in the canonical space and resulting in smoother surfaces.
+Furthermore, we introduce a filtering method to remove geometry floaters by filtering the depth image using the rendered high-quality RGB mask.
 This approach allows for an accurate surface mesh to be obtained via the TSDF method.
@@
 Our framework takes into account both the dynamic and geometric properties of objects.
-  • To remove Geometry floaters, we propose to filter the depth image using a mask extracted from the rendered high-quality RGB image.
-  • Experiments demonstrate that our D-2DGS achieves state-of-the-art (SOTA) reconstruction quality compared with other advanced representations.
-Figure 1: Illustration of Geometry floaters.
-In the canonical space, all the spatial Gaussian primitives are connected by the control points.
+  • To extract a clean mesh, it is necessary to remove the Geometry floaters generated during reconstruction.
+Therefore, we propose to filter the depth image using a mask extracted from the rendered high-quality RGB image.
+  • Experiments demonstrate that our D-2DGS achieves state-of-the-art (SOTA) reconstruction/mesh rendering quality compared with other advanced representations.
+Figure 2.
+Illustration of Geometry floaters.
 After deformation, both objects Gaussians and floater Gaussians are transformed by control points.
 The similar color to the background makes it hard to be pruned, which causes inaccurate geometry.
-Related Works
-3D/4D Representations
+2. Related Works
+2.1. 3D/4D Representations
 Representing the 4D world and rendering photo-realistic novel view images has been an important and challenging topic.
-Explicit representations such as mesh (Collet et al.
-2015; Li et al.
-2018; Guo et al.
-2015; Su et al.
-2020), voxels (Guo et al.
-2019; Hu et al.
-2022; Li et al.
-2017), and point clouds enjoy editable manipulations and GPU-friendly applications but suffer from strict multi-view constraints or more sensors.
-Recently, implicit (Martin-Brualla et al.
-2021; Park et al.
-2021a, b) representations have demonstrated their efficiency for novel view synthesis from sparse input and many approaches (Xu et al.
-2022; Fridovich-Keil et al.
-2023; Cao and Johnson 2023; Wang et al.
-2023a; Shao et al.
-2023; Gan et al.
-2023; Li et al.
-2022) combine the benefits of these representations.
-Notably, Gaussian Splatting (Kerbl et al.
-2023; Yang et al.
-2023b; Wu et al.
-2024; Yang et al.
-2023a; Zhang et al.
-2024) (GS) based representations as an efficient mixture presentation, show the ability to render high-quality novel views and support many downstream tasks (Jiang et al.
-2024; Ji et al.
-2024; Yi et al.
-2024a, b; Yin et al.
-2023; Liang et al.
-2024).
-Recent 4D representation, SC-GS (Huang et al.
-2024b), proposes a sparse control points method to control the deformation of 3D Gaussians, which reaches outstanding NVS quality and high training efficiency.
-However, most 3D/4D representations are only for NVS or extract mesh from static scenes and there is less research focus on extracting accurate mesh from dynamic objects in monocular input.
-Our D-2DGS shows the ability to reconstruct the accurate 3D mesh from 2D images captured in the 4D real world with high efficiency.
-Geometry Reconstruction
-Geometry reconstructions are mainly from streamable multiview (Collet et al.
-2015; Li et al.
-2018; Guo et al.
-2015; Su et al.
-2020) constraints.
-Recently, 3DGS (Kerbl et al.
-2023) is proven to enjoy high training efficiency and shows great potential to reconstruct high-quality mesh from spatial Gaussians.
+Explicit representations such as mesh (Li et al., 2018; Guo et al., 2015), voxels (Guo et al., 2019; Hu et al., 2022; Li et al., 2017), and point clouds (Xu et al., 2022) enjoy editable manipulations and GPU-friendly applications but suffer from strict multi-view constraints or more sensors.
+Recently, implicit (Martin-Brualla et al., 2021; Park et al., 2021a, b) representations have demonstrated their efficiency for novel view synthesis from sparse input, and many approaches (Fridovich-Keil et al., 2023; Cao and Johnson, 2023; Shao et al., 2023; Fang et al., 2022) combine the benefits of these representations.
+Notably, Gaussian Splatting (Kerbl et al., 2023; Yang et al., 2024b; Wu et al., 2024; Yang et al., 2024a; Zhang et al., 2024; Lei et al., 2025) (GS) based representations as an efficient mixture presentation, show the ability to render high-quality novel views and support many downstream tasks (Jiang et al., 2024; Ji et al., 2024; Yi et al., 2024a, b).
+Recent 4D representation, SC-GS (Huang et al., 2024a), proposes a sparse control points method to control the deformation of 3D Gaussians, which reaches outstanding NVS quality and high training efficiency.
+However, most 3D/4D representations are only for NVS or extracting mesh from static scenes, and there is less research focus on extracting an accurate mesh from dynamic objects.
+Our D-2DGS shows the ability to reconstruct an accurate 3D mesh from 2D images captured in the 4D world.
+2.2. Geometry Reconstruction
+Geometry reconstructions are mainly from streamable multiview constraints (Collet et al., 2015; Li et al., 2018; Guo et al., 2015).
+Recently, 3DGS (Kerbl et al., 2023) is proven to enjoy high training efficiency and shows great potential to reconstruct high-quality mesh from spatial Gaussians.
 Many works are proposed to improve the geometric accuracy of 3DGS.
-SuGaR (Guédon and Lepetit 2024) introduces a regularization term that encourages the Gaussians to align well with the surface of the scene.
-NeuSG (Chen, Li, and Lee 2023) jointly optimizes 3DGS and neural implicit models, combining the advantages of both, enabling the generation of complete surfaces with high details.
-Similarly, 3DGSR (Lyu et al.
-2024) also encourages the alignment of 3D Gaussian and SDF to improve the surface reconstruction effect.
-2D Gaussian Splatting (2DGS) (Huang et al.
-2024a)is a breakthrough in mesh reconstruction since the view geometry consistency is maintained by planar Gaussian disks.
+SuGaR (Guédon and Lepetit, 2024) introduces a regularization term that encourages the Gaussians to align well with the surface of the scene.
+NeuSG (Chen et al., 2023) jointly optimizes 3DGS and neural implicit models, combining the advantages of both, enabling the generation of complete surfaces with high details.
+Similarly, 3DGSR (Lyu et al., 2024) also encourages the alignment of 3D Gaussian and SDF to improve the surface reconstruction effect.
+2D Gaussian Splatting (2DGS) (Huang et al., 2024b)is a breakthrough in mesh reconstruction since the view geometry consistency is maintained by planar Gaussian disks.
+Recently, 3D foundation models  (Xiang et al., 2025; Tochilkin et al., 2024; Li et al., 2025; Ye et al., 2025) enable high-quality mesh generation from sparse input.
 The above geometric reconstruction methods are all for static scenes.
 Recently, some works have been attempted on the geometric reconstruction of dynamic scenes.
-Dynamic Gaussian Mesh (DG-Mesh) (Liu, Su, and Wang 2024) aims to reconstruct a high-fidelity and temporally consistent mesh from a given monocular video.
-DG-Mesh introduces Gaussian grid anchoring to encourage uniformly distributed Gaussian distributions and obtains better grid reconstruction by grid densification and pruning of deformed Gaussian distributions.
-MaGS (Ma, Luo, and Yang 2024) propose mutual-adsorbed mesh Gaussian representation to refine the generated mesh.
-Vidu4D (Wang et al.
-2024b) uses Dynamic Gaussian Surfels as scene representation primitives, achieving text to 4D generation with high-quality appearance and geometry.
+Dynamic Gaussian Mesh (DGMesh) (Liu et al., 2025) aims to reconstruct a high-fidelity and temporally consistent mesh from a given monocular video.
+DGMesh introduces Gaussian grid anchoring to encourage uniformly distributed Gaussian distributions and obtains better grid reconstruction by grid densification and pruning of deformed Gaussian distributions.
+MaGS (Ma et al., 2024) proposes a mutual-adsorbed mesh Gaussian representation to refine the generated mesh.
+Vidu4D (Wang et al., 2024b) uses Dynamic Gaussian Surfels as scene representation primitives, achieving text-to-4D generation with high-quality appearance and geometry.
+Here are many concurrent works (Ma et al., 2024; Li et al., 2024) that propose to use a 3D Gaussian representation to reconstruct a mesh from dynamic scenes.
+Papers(Cai et al., 2024; Wang et al., 2024a) adopt a deformation field network and Gaussian surfels to model the scene’s deformation.
 Our methods propose to model Gaussian surfels’ motion by sparse-controlled points, maintaining consistent and smooth surface mesh generation.
-Preliminaries
-2D Gaussian Splatting
+3. Preliminaries
+3.1. 2D Gaussian Splatting
 2D Gaussian Splatting (2DGS) simplifies 3D modeling through 2D Gaussians.
@@
 The 2D Gaussian function can be parameterized on the local tangent plane in world space as:
-$P(u,v)=P_{k}+s_{u}\mathbf{t}_{u}u+s_{v}\mathbf{t}_{v}v.$ (1)
+(1) $P(u,v)=P_{k}+s_{u}\mathbf{t}_{u}u+s_{v}\mathbf{t}_{v}v.$
 For a point $\mathbf{u}=(u,v)$ in the $uv$ space, its 2D Gaussian value is:
-$\mathcal{G}(\mathbf{u})=exp(-\frac{u^{2}+v^{2}}{2}).$ (2)
+(2) $\mathcal{G}(\mathbf{u})=exp(-\frac{u^{2}+v^{2}}{2}).$
 The parameters of each 2D Gaussian also include opacity $\alpha$ and view-dependent appearance $c$ parameterized with spherical harmonics.
 2D Gaussians are sorted by their center depth and organized into tiles based on their bounding boxes. The alpha-weighted appearance is integrated from front to back using volumetric alpha blending:
-$c(x)=\sum_{i=1}\mathbf{c}_{i}\alpha\mathcal{G}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-\alpha_{j}\mathcal{G}(\mathbf{u}(x))).$ (3)
-However, 2D Gaussians can only reconstruct 3D mesh from static scenes.
-We propose to use Sparse-Controlled points to model 2D Gaussians deformation and extend 2D Gaussians representing the 4D world successfully.
-Sparse-Controlled Gaussian Splatting
-Sparse-Controlled Gaussian Splatting (SC-GS) (Huang et al.
-2024b) shows excellent performance in the task of synthesizing novel views in dynamic scenes.
-It uses sparse control points to control dense 3D Gaussians in the scene.
-For each control point at timestep $t$, SC-GS uses MLP $\Phi$ to predict a rotation matrix $R_{i}^{t}\in\mathbf{SE}(3)$ and rotation matrix $T_{i}^{t}\in\mathbb{R}^{3}$.
-For each Gaussian $G_{j}:(\mu_{j},q_{j},s_{j},\sigma_{j},sh_{j})$, use k-nearest neighbor (KNN) search to get the $K$ neighboring control points $\{p_{i},i=0,1,...,k\}$.
-SC-GS employs LBS (Sumner, Schmid, and Pauly 2007) $\mathcal{I}$ to compute the warped Gaussian $\mu_{j}^{t}$ and $q_{j}^{t}$:
-$\mu_{j}^{t}=\sum_{k\in\mathcal{N}_{j}}w_{jk}(R_{k}^{t}(\mu_{j}-p_{k}))+p_{k}+T_{k}^{t}),$ (4)
-$q_{j}^{t}=(\sum_{k\in\mathcal{N}_{j}}w_{jk}r_{k}^{t})\otimes q_{j},$ (5)
-where $R_{k}^{t}$ and $r_{k}^{t}$ are the matrix and quaternion representations of the predicted rotation at control point $k$, respectively. The weight $w_{jk}$ is calculated by:
-$w_{jk}=\frac{\hat{w}_{jk}}{\sum_{k\in\mathcal{N}_{j}}\hat{w}_{jk}},\mathrm{where}~{}\hat{w}_{jk}=exp(-\frac{d_{jk}^{2}}{2o_{k}^{2}}),$ (6)
-where $d_{jk}$ is the distance between the Gaussian and the adjacent control point, and $o_{k}$ is the learned radius parameter of the control point.
-However, SC-GS (Huang et al.
-2024b) mainly chooses 3D-GS (Kerbl et al.
-2023) as its geometry representation, leading to multiview inconsistency.
-Meanwhile, There are fewer geometry constraints during the optimization of canonical Gaussians, we mainly choose 2DGS (Huang et al.
-2024a) as basic geometry, and propose an outlier removal method to overcome the Geometry floaters as shown in Fig.
-1 from the sparse monocular input.
-Method
-Figure 2: Framework of our D-2DGS.
+(3) $c(x)=\sum_{i=1}\mathbf{c}_{i}\alpha\mathcal{G}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-\alpha_{j}\mathcal{G}(\mathbf{u}(x))).$
+However, 2D Gaussians can only reconstruct a 3D mesh from static scenes.
+We propose to use Sparse-Controlled points to model 2D Gaussians deformation and extend 2D Gaussians to represent the 4D world successfully.
+Figure 3.
+Framework of our D-2DGS.
 Sparse points are bonded with canonical 2D Gaussians.
@@
 To get high-quality meshes, depth images are filtered by rendered images with RGB mask, and then TSDF is applied on multiview depth images and RGB images.
-Dynamic 2D Gaussians Framework
+3.2. Sparse-Controlled Gaussian Splatting
+Sparse-Controlled Gaussian Splatting (SC-GS) (Huang et al., 2024a) shows excellent performance in the task of synthesizing novel views in dynamic scenes.
+It uses sparse control points to control dense 3D Gaussians in the scene.
+For each control point at timestep $t$, SC-GS uses MLP $\Phi$ to predict a rotation matrix $R_{i}^{t}\in\mathbf{SE}(3)$ and translation matrix $T_{i}^{t}\in\mathbb{R}^{3}$.
+For each Gaussian $G_{j}:(\mu_{j},q_{j},s_{j},\sigma_{j},sh_{j})$, use k-nearest neighbor (KNN) search to get the $K$ neighboring control points $\{p_{i},i=0,1,...,k\}$.
+SC-GS employs LBS (Sumner et al., 2007) $\mathcal{I}$ to compute the warped Gaussian $\mu_{j}^{t}$ and $q_{j}^{t}$:
+(4) $\mu_{j}^{t}=\sum_{k\in\mathcal{N}_{j}}w_{jk}(R_{k}^{t}(\mu_{j}-p_{k}))+p_{k}+T_{k}^{t}),$
+(5) $q_{j}^{t}=(\sum_{k\in\mathcal{N}_{j}}w_{jk}r_{k}^{t})\otimes q_{j},$
+where $R_{k}^{t}$ and $r_{k}^{t}$ are the matrix and quaternion representations of the predicted rotation at control point $k$, respectively. The weight $w_{jk}$ is calculated by:
+(6) $w_{jk}=\frac{\hat{w}_{jk}}{\sum_{k\in\mathcal{N}_{j}}\hat{w}_{jk}},\mathrm{where}~\hat{w}_{jk}=exp(-\frac{d_{jk}^{2}}{2o_{k}^{2}}),$
+where $d_{jk}$ is the distance between the Gaussian and the adjacent control point, and $o_{k}$ is the learned radius parameter of the control point.
+However, SC-GS (Huang et al., 2024a) mainly chooses 3D-GS (Kerbl et al., 2023) as its geometry representation, leading to multiview inconsistency.
+Meanwhile, there are fewer geometry constraints during the optimization of canonical Gaussians, we mainly choose 2DGS (Huang et al., 2024b) as basic geometry, and propose an outlier removal method to overcome the Geometry floaters as shown in Fig.
+2.
+4. Method
+4.1. Dynamic 2D Gaussians Framework
 The pipeline of our framework is illustrated in Fig.
-2.
+3.
 We adopt 2D Gaussian primitives $\mathcal{G}$ for scene representation.
-We follow deformation-based Gaussian splatting to compute the deformation of each 2D Gaussians at timestamp $t_{j}$:
-$\mathcal{G}^{j}=Deform(\mathcal{G},t_{j},\mathbf{P};\Phi),$ (7)
-where $Deform$ denotes deformation process, $\mathbf{P}=\{p_{i}\}_{i=1}^{N}$ denotes sparse control points.
-Then $\mathcal{G}^{j}$ are used to render the RGB image $\mathbf{I}$ and the depth image $\mathbf{D}$ by differential splatting algorithm (Huang et al.
-2024b; Yifan et al.
-2019) $\mathcal{S}$ with a camera parameters $[\mathbf{R},\mathbf{t}]$:
-$\mathbf{I},\mathbf{D}=\mathcal{S}\left(\mathcal{G}^{j};\theta),[\mathbf{R},\mathbf{t}]\right).$ (8)
+We follow deformation-based Gaussian splatting to compute the deformation of each 2D Gaussian at timestamp $t_{j}$:
+(7) $\mathcal{G}^{j}=\Phi(\mathcal{G},t_{j},\mathbf{P}),$
+where $\Phi$ denotes deformation network, $\mathbf{P}=\{p_{i}\}_{i=1}^{N}$ denotes sparse control points.
+Then $\mathcal{G}^{j}$ are used to render the RGB image $\mathbf{I}$ and the depth image $\mathbf{D}$ by differential splatting algorithm (Huang et al., 2024a; Yifan et al., 2019) $\mathcal{S}$ with a camera parameters $\mathbf{R},\mathbf{t}]$:
+(8) $\mathbf{I},\mathbf{D}=\mathcal{S}\left((\mathcal{G}^{j};\theta),[\mathbf{R},\mathbf{t}]\right).$
 We then extract a mask $\mathbf{M}$ from the rendered image $\mathbf{I}$ and apply it to the depth image $\mathbf{D}$ to filter out floaters:
-$\mathbf{D}^{\prime}=Flitering(\mathbf{I},\mathbf{D}),$ (9)
+(9) $\mathbf{D}^{\prime}=Flitering(\mathbf{I},\mathbf{D}),$
 the $Flitering$ progress will be discussed later.
 Finally, the refined depth image $\mathbf{D}^{\prime}$ is used to extract the mesh using the Truncated Signed Distance Function (TSDF).
-Sparse point controlled 2D Gaussians
+4.2. Sparse point controlled 2D Gaussians
 To accurately extract dynamic mesh sequences, it is essential to account for both the dynamic and geometric characteristics of the modeled objects.
 We leverage the multi-view geometric consistency of 2D Gaussian representations, selecting 2D Gaussians $\mathcal{G}:(\mu,q,s,\sigma,\textit{sh})$ as the foundational primitive for scene representation.
-The accuracy of motion is ensured by sparse control points which are proposed in SC-GS (Huang et al.
-2024b).
+The accuracy of motion is ensured by sparse control points, which are proposed in SC-GS (Huang et al., 2024a).
 The sparse controlled points $\mathbf{P}$ and deformation network $\Phi(p_{i};\theta)$ are proposed to model the accurate deformation.
@@
 The control signals at $t_{j}$ are computed by the MLP $\Phi$:
-$[R_{i}^{j},T_{i}^{j}]=\Phi(p_{i},t_{j}).$ (10)
-Given $K$ adjacent control points $\mathbf{p}=\{p_{0},\dots,p_{k}\}$, the 6-DoF transform $\{R^{j},T^{j}\}$ can be computed by a deformation network $\Phi$ applying on $p$, and the deformed 2D Gaussians $\mathcal{G}^{j}$ at $t_{j}$ can be interpolated using the LBS (Sumner, Schmid, and Pauly 2007) $\mathcal{I}$ including Eq.
+(10) $R_{i}^{j},T_{i}^{j}]=\Phi(p_{i},t_{j}).$
+Given $K$ adjacent control points $\mathbf{p}=\{p_{0},\dots,p_{k}\}$, the 6-DoF transform $\{R^{j},T^{j}\}$ can be computed by a deformation network $\Phi$ applying on $p$, and the deformed 2D Gaussians $\mathcal{G}^{j}$ at $t_{j}$ can be interpolated using the LBS (Sumner et al., 2007) $\mathcal{I}$ including Eq.
 4 and Eq.
 5 :
-$\mathcal{G}^{j}=\mathcal{I}\left(\{R^{j},T^{j}\},\mathcal{G}\right).$ (11)
+(11) $\mathcal{G}^{j}=\mathcal{I}\left(\{R^{j},T^{j}\},\mathcal{G}\right).$
 This approach ensures that both the temporal dynamics and geometric consistency are preserved, facilitating the extraction of high-fidelity dynamic mesh sequences.
-Figure 3: Mesh and mesh rendering visualization results of the D-NeRF dataset.
-Method Lego Bouncingballs Jumpingjacks Hook
-[…]
-DG-mesh* 23.474 0.903 0.103 28.045 0.972 0.088 31.667 0.977 0.036 29.217 0.965 0.049
-D-3DGS 22.534 0.884 0.115 23.667 0.956 0.102 28.878 0.972 0.039 27.499 0.959 0.048
-SCGS 22.784 0.882 0.113 25.231 0.956 0.103 29.534 0.976 0.035 27.095 0.961 0.048
-D-2DGS (Ours) 23.293 0.887 0.112 27.786 0.969 0.073 29.293 0.974 0.032 27.802 0.962 0.042
-Method Mutant Standup Trex Hellwarrior
-[…]
-DG-mesh* 31.369 0.970 0.043 32.220 0.979 0.036 29.559 0.970 0.044 27.602 0.971 0.055
-D-3DGS 28.342 0.957 0.048 29.946 0.977 0.031 28.061 0.958 0.054 25.986 0.965 0.053
-SCGS 28.002 0.955 0.047 29.273 0.977 0.032 28.521 0.962 0.050 26.419 0.967 0.048
-D-2DGS (Ours) 28.120 0.960 0.042 29.512 0.976 0.028 28.677 0.967 0.043 25.441 0.958 0.051
-Table 1: Rendering quality of extracted meshes on the D-NeRF (Pumarola et al.
-2021) dataset.
-The color of each cell means: best, second best.
-(*DG-mesh needs to query additional MLPs when rendering mesh.)
-Depth filtering
+Figure 4. Mesh and mesh rendering visualization results of the D-NeRF dataset.
+4.3. Depth filtering
 During the optimization process, Geometry floaters manifest as numerous background-colored artifacts, despite the high quality of the rendered image.
@@
 The extraction of the mask $\mathbf{M}$ from the rendered image $\mathbf{I}$ can be mathematically described as follows:
-$\mathbf{M}(x)=\mathbf{1}[\mathbf{I}(x)=bg],$ (12)
+(12) $\mathbf{M}(x)=\mathbf{1}[\mathbf{I}(x)=bg],$
 where $\mathbf{1}[\cdot]$ is the indicator function that returns 1 if the color of $\mathbf{I}(x)$ is different from the background color $bg$, and 0 otherwise.
 The mask $\mathbf{M}$ is then used to refine the depth image $\mathbf{D}$ by element-wise multiplication:
-$\mathbf{D}^{\prime}=\mathbf{D}\odot\mathbf{M},$ (13)
+(13) $\mathbf{D}^{\prime}=\mathbf{D}\odot\mathbf{M},$
 where $\odot$ represents the element-wise product.
 The resulting filtered depth image $\mathbf{D}^{\prime}$ effectively removes the floating noise points, leading to a cleaner mesh extraction with reduced artifacts.
-Loss Function
+4.4. Loss Function
 In order to normalize the spatial distribution of Gaussians, we introduce the Depth Distortion and Normal Consistency regularization terms in 2DGS.
 The Depth Distortion regularization term encourages Gaussians to be distributed at the same depth as much as possible, as shown in the following formula:
-$L_{d}=\sum_{i,j}\omega_{i}\omega_{j}\left|z_{i}-z_{j}\right|,$ (14)
+(14) $L_{d}=\sum_{i,j}\omega_{i}\omega_{j}\left|z_{i}-z_{j}\right|,$
 where $i$ indexes over intersected splats along the ray, $\omega_{i}$ denotes the blending weight of the intersection point.
-$\omega_{i}=\alpha_{i}\mathcal{G}_{i}(u(x)){\textstyle\prod_{j=1}^{i-1}}(1-\alpha\mathcal{G}_{j}(u(x)))$ $z_{i}$ is the depth of the intersection points.
+$\omega_{i}=\alpha_{i}\mathcal{G}_{i}(u(x)){\textstyle\prod_{j=1}^{i-1}}(1-\alpha\mathcal{G}_{j}(u(x)))$ .
+$z_{i}$ is the depth of the intersection points.
 The Normal Consistency regularization term can make the normals of 2D Gaussians as close as possible to the normals of the object:
-$L_{n}=\sum_{i}\omega_{i}(1-n_{i}^{\mathrm{T}}N),$ (15)
+(15) $L_{n}=\sum_{i}\omega_{i}(1-n_{i}^{\mathrm{T}}N),$
 where $n_{i}$ represents the normal of Gaussians, $N$ normal estimated by the nearby depth point $p$, it is calculated by the following formula:
-$N(x,y)=\frac{\nabla_{x}p\times\nabla_{y}p}{\left|\nabla_{x}p\times\nabla_{y}p\right|}.$ (16)
+(16) $N(x,y)=\frac{\nabla_{x}p\times\nabla_{y}p}{\left|\nabla_{x}p\times\nabla_{y}p\right|}.$
 Therefore, the overall loss function is:
-$L_{all}=L_{1}(I_{r},I_{g})+\lambda_{s}L_{ssim}(I_{r},I_{g})+\lambda_{n}L_{n}+\lambda_{d}L_{d},$ (17)
+(17) $L_{all}=L_{1}(I_{r},I_{g})+\lambda_{s}L_{ssim}(I_{r},I_{g})+\lambda_{n}L_{n}+\lambda_{d}L_{d},$
 where $I_{r}$ represents the rendered image, and $I_{g}$ represents the Ground Truth. $\lambda_{s}$ and $\lambda_{n}$ are the weights of each loss function respectively.
-Method D-NeRF datasets DG-Mesh datasets
-[…]
-DG-Mesh* 29.114 0.963 0.057 0.698 0.123
-D-3DGS 26.864 0.954 0.061 0.895 0.115
-SCGS 27.107 0.954 0.060 0.856 0.115
-D-2DGS(ours) 27.490 0.957 0.052 0.851 0.132
-Table 2: Average results on both D-NeRF datasets and DG-Mesh datasets. (*DG-mesh needs to query additional MLPs when rendering mesh.)
-Mesh Extraction
-The mesh at any normalized timestamp $t_{i}\in[0,1]$ can be extracted from the trained model.
-Given $t_{i}$, the model renders multi-view images $\mathbf{I}(t_{i})$ and depth images $\mathbf{D}(t_{i})$.
-The mesh is then obtained by applying TSDF using Open3D (Zhou, Park, and Koltun 2018):
-$\text{Mesh}=\text{TSDF}\left(\{\mathbf{I}_{i},\mathbf{D}_{i}^{\prime}\mid i\in T\}\right).$ (18)
-Experiment
-Experimental Setups
-We mainly compare some novel view synthesis methods for dynamic scenes based on Gaussian Splatting, including DG-Mesh (Liu, Su, and Wang 2024), Deformable 3DGS (Yang et al.
-2023a), and SCGS (Huang et al.
-2024b).
-We evaluate the quality of meshes extracted by these methods on the D-NeRF (Pumarola et al.
-2021) and DG-Mesh (Liu, Su, and Wang 2024) datasets.
-Both datasets have large and non-rigid motions and belong to efficient multiview setups as referred to the Dycheck (Gao et al.
-2022).
-Each timestamp has only one view which is randomly selected.
-Among them, the D-NeRF dataset does not contain the ground truth of the mesh but the DG-Mesh dataset includes the GT mesh.
-Therefore, We report the PSNR, LPIPS, and SSIM to evaluate the rendered images in the D-NeRF datasets and use Chamfer Distance (CD) and Earth Mover’s Distance (EMD) as mesh evaluation metrics.
-The mesh extracted by deformable 3DGS and SC-GS following the same settings as ours and 2D-GS (Huang et al.
-2024a), and it is worth noting that when DG-Mesh renders an image, two additional MLPs should be maintained when querying the color of the mesh.
-Figure 4: Mesh and mesh rendering visualization results of the DG-Mesh dataset.
-Implementation Details
-We build dynamic 2DGS upon the open-source SCGS (Huang et al.
-2024b) codebase.
+4.5. Mesh Extraction
+The mesh at any normalized timestamp $t\in[0,1]$ can be extracted from the trained model.
+Given $t$, the model renders multi-view images $\mathbf{I}(t)$ and filtered depth images $\mathbf{D}^{\prime}(t)$.
+The mesh is then obtained by applying TSDF (Zhou et al., 2018):
+(18) $\text{Mesh}(t)=\text{TSDF}\left(\mathbf{I}(t),\mathbf{D}^{\prime}(t)\right).$
+The mesh of an object at any time can be extracted by changing the $t$ value.
+4.6. Post-processing
+Experiments have found that the extracted mesh occasionally has problems with damage and holes, so we use MeshLab to repair small holes in the extracted mesh.
+The hole-filling algorithm first detects the holes on the mesh, then sets the maximum hole size allowed to be filled, and finally generates new triangular patches to close the holes for the hole boundaries that meet the conditions.
+5. Experiment
+5.1. Experimental Setups
+We mainly compare some novel view synthesis methods for dynamic scenes, including TiNeuVox-B (Fang et al., 2022), DGMesh (Liu et al., 2025), DynaSurfGS (Cai et al., 2024), Deformable 3DGS (Yang et al., 2024a), and SCGS (Huang et al., 2024a).
+We evaluate the quality of meshes extracted by these methods on the D-NeRF (Pumarola et al., 2021) and DGMesh (Liu et al., 2025) datasets.
+Both datasets have large and non-rigid motions and belong to efficient multiview setups as referred to the Dycheck (Gao et al., 2022).
+Each timestamp has only one view, which is randomly selected.
+Among them, the D-NeRF dataset does not contain the ground truth of the mesh.
+Therefore, we report the PSNR, LPIPS, and SSIM of the rendered images by mesh to evaluate the mesh quality in the D-NeRF datasets.
+For the DGMesh dataset, we report the CD and EMD to evaluate the quality of the generated mesh.
+It is worth noting that when DGMesh renders an image, two additional MLPs should be maintained when querying the color of the mesh.
+In addition, DGMesh uses a differentiable Poisson Solver and differentiable Marching Cubes to extract the mesh.
+Other methods use TSDF to extract mesh like 2D-GS (Huang et al., 2024b).
+5.2. Implementation Details
+We build dynamic 2DGS upon the open-source SCGS (Huang et al., 2024a) codebase.
 Training iterations for all scenarios are set to 80,000.
@@
 When the mesh is extracted, the original dynamic 2DGS representations can be removed.
-Method GS render image Mesh
-Color Smooth Detail
-DG-Mesh
-D-3DGS
-SCGS
-Ours
-Table 3: The comparison of different models.
+The indicators of TiNeuVox-B and DGMesh come from the DGMesh paper.
+Method Lego Bouncingballs Jumpingjacks Hook
 […]
-w/o Depth filtering 0.425 0.110
-w/o Normal Consistency 0.443 0.101
-w/o Depth Distortion 0.350 0.109
-Full model 0.328 0.110
-Table 4: Ablation experiment results of the “bird” sample in the DG-Mesh dataset.
-Results
-The results on the D-NeRF (Pumarola et al.
-2021) dataset are shown in Fig.
-3.
+TiNeuVox-B (Fang et al., 2022) 21.927 0.843 0.126 24.819 0.947 0.101 23.621 0.932 0.075 21.429 0.908 0.085
+D-3DGS (Yang et al., 2024a) 22.534 0.884 0.115 23.667 0.956 0.102 28.878 0.972 0.039 27.499 0.959 0.048
+SCGS (Huang et al., 2024a) 22.784 0.882 0.113 25.231 0.956 0.103 29.534 0.976 0.035 27.095 0.961 0.048
+DGMesh* (Liu et al., 2025) 21.289 0.838 0.159 29.145 0.972 0.099 31.769 0.977 0.045 27.884 0.954 0.074
+D-2DGS (Ours) 23.293 0.887 0.112 27.786 0.969 0.073 29.293 0.974 0.032 27.802 0.962 0.042
+Method Mutant Standup Trex Hellwarrior
+[…]
+TiNeuVox-B (Fang et al., 2022) 22.967 0.925 0.064 24.263 0.941 0.051 24.219 0.927 0.070 18.657 0.917 0.118
+D-3DGS (Yang et al., 2024a) 28.342 0.957 0.048 29.946 0.977 0.031 28.061 0.958 0.054 25.986 0.965 0.053
+SCGS (Huang et al., 2024a) 28.002 0.955 0.047 29.273 0.977 0.032 28.521 0.962 0.050 26.419 0.967 0.048
+DGMesh* (Liu et al., 2025) 30.400 0.968 0.055 30.208 0.974 0.051 28.951 0.959 0.065 25.460 0.959 0.084
+D-2DGS (Ours) 28.120 0.960 0.042 29.512 0.977 0.028 28.677 0.967 0.043 25.498 0.958 0.049
+Table 1.
+Rendering quality of extracted meshes on the D-NeRF (Pumarola et al., 2021) dataset (Background: white).
+The color of each cell means: best, second best.
+(*DGMesh uses differentiable Marching Cubes to extract mesh, and the extracted mesh lacks details.)
+Figure 5. Mesh visualization results of the DGMesh dataset.
+Method D-NeRF datasets DGMesh datasets
+[…]
+TiNeuVox-B (Fang et al., 2022) 22.738 0.918 0.086 1.839 0.130
+D-3DGS (Yang et al., 2024a) 26.864 0.954 0.061 0.895 0.115
+SCGS (Huang et al., 2024a) 27.107 0.954 0.060 0.856 0.115
+DGMesh* (Liu et al., 2025) 28.137 0.950 0.057 0.698 0.123
+DynaSurfGS (Cai et al., 2024) - 0.910 0.123
+D-2DGS(Ours) 27.498 0.957 0.052 0.647 0.112
+Table 2. Average results on both D-NeRF datasets and DGMesh dataset.
+5.3. Results
+The results on the D-NeRF (Pumarola et al., 2021) dataset are shown in Fig.
+4.
 The ideal surface should be as smooth as possible while still containing the details of the object.
-The mesh extracted by DG-Mesh (Liu, Su, and Wang 2024) has a low resolution and lacks some details of objects, such as the hands of the ‘jumpingjacks’.
+The mesh extracted by DGMesh (Liu et al., 2025) has a low resolution and lacks some details of objects, such as the hands of the ‘jumpingjacks’.
 Deformable-3DGS and SCGS both use 3D Gaussians as the basic representational primitives, which lack geometric accuracy.
@@
 The three comparative methods are unable to simulate smooth planar and spherical surfaces, as demonstrated by the “bouncingballs” example.
-The sphere appears broken and the chassis is uneven due to the influence of lighting.
+The sphere appears broken, and the chassis is uneven due to the influence of lighting.
 In comparison, our Dynamic 2DGS highlights reconstructing rich details and modeling smooth surfaces.
@@
 1.
-When DG-Mesh renders an image, it needs to query additional MLPs to obtain the color of the mesh, thus achieving higher PSNR and SSIM.
-The proposed D-2DGS achieved higher indicators compared with D-3DGS and SCGS.
+When DGMesh renders an image, it needs to query additional MLPs to obtain the color of the mesh, thus achieving higher PSNR.
+The proposed D-2DGS achieved higher indicators.
+The average metrics are shown in Tab.
+2 and Tab.
+4.
 As shown in Fig.
-4, on the DG-Mesh dataset, the proposed method can extract smoother meshes compared to D-3DGS and SCGS.
-At the same time, compared to DG-Mesh, the proposed method can show the details of objects, such as the mouth of the ‘horse’.
-The average metrics of the D-NeRF and DG-Mesh datasets are shown in Tab.
+5, on the DGMesh dataset, the proposed method can extract smoother meshes compared to D-3DGS and SCGS.
+Compared to DGMesh, the proposed method can show the details of objects, such as the mouth of the ‘horse’.
+The indicators on the DGMesh dataset are shown in Tab.
+3 and Tab.
 2.
-The data in the DG-Mesh dataset has less details, so the proposed model is comparable to other models in terms of metrics.
-We summarize the methods on the Tab.
-3.
-DG-Mesh produces low-quality rendered images through Gaussian rendering, and the extracted meshes rely on MLP modeling for color, and lacking in detail.
-Meanwhile, D-3DGS and SCGS extract meshes that are over-fitted to details, resulting in surfaces that are not sufficiently smooth.
-Figure 5: Ablation Study on Sparse point control.
-Ablation Study
+The proposed method achieves the lowest CD and EMD value, demonstrating the effectiveness of the proposed method on this dataset.
+We evaluate the quality of the mesh from three aspects: whether it is clean (i.e., whether there are floating objects), whether it is smooth, and whether it has details.
+The image quality rendered by the DGMesh method through Gaussian is relatively low.
+In addition, the mesh extracted by DGMesh lacks details.
+The mesh extracted by D-3DGS and SCGS is prone to floaters.
+In addition, its surface is uneven, prone to potholes, and not smooth enough.
+In comparison, the mesh extracted by our method not only has better details but is also relatively smooth.
+Method Duck Horse Bird
+[…]
+TiNeuVox-B (Fang et al., 2022) 0.969 0.059 1.918 0.246 8.264 0.215
+D-3DGS (Yang et al., 2024a) 1.064 0.076 0.298 0.114 0.427 0.123
+SCGS (Huang et al., 2024a) 1.001 0.074 0.232 0.118 0.336 0.121
+DGMesh (Liu et al., 2025) 0.782 0.048 0.299 0.168 0.557 0.128
+D-2DGS (Ours) 0.841 0.057 0.327 0.134 0.328 0.110
+Method Beagle Torus2sphere Girlwalk
+[…]
+TiNeuVox-B (Fang et al., 2022) 0.874 0.129 2.115 0.203 0.568 0.184
+D-3DGS (Yang et al., 2024a) 0.585 0.101 2.655 0.155 0.340 0.122
+SCGS (Huang et al., 2024a) 0.528 0.100 2.716 0.155 0.321 0.124
+DGMesh (Liu et al., 2025) 0.626 0.114 1.515 0.126 0.406 0.153
+D-2DGS (Ours) 0.553 0.115 1.502 0.139 0.332 0.116
+Table 3. Mesh reconstruction quality comparison results of DGMesh dataset.
+Method Train Time Render Speed Memory Storage
+TiNeuVox-B (Fang et al., 2022) 28 mins 1.5 FPS 5574 MB 48 MB
+D-3DGS (Yang et al., 2024a) 10 mins 190 FPS 4270 MB 10 MB
+SCGS (Huang et al., 2024a) 16 mins 215 FPS 3200 MB 26 MB
+DGMesh (Liu et al., 2025) 90 mins 140 FPS 9425 MB 17 MB
+D-2DGS (Ours) 32 mins 71 FPS 3690 MB 20 MB
+Table 4. Comprehensive comparison of different methods.
+Figure 6. Ablation Study on Sparse point control.
+5.4. Ablation Study
 Sparse point control.
@@
 As shown in Fig.
-5, in areas with significant motion, such as a person’s hand and a horse’s hoof, sparse point control allows for the extraction of a more detailed mesh.
-Figure 6: Ablation Study on Depth filtering.
+6, in areas with significant motion, such as a person’s hand and a horse’s hoof, sparse point control allows for the extraction of a more detailed mesh.
+This shows that sparse point control can model finer motions, thus helping to extract more detailed meshes.
+Figure 7. Ablation Study on Depth filtering.
 Depth filtering.
@@
 After extracting the mesh through TSDF fusion, some floating objects will appear around the mesh of the object.
-With the help of high-quality renderings, the object mask is extracted and the depth image is masked.
-The noise in the depth image can be filtered out, thereby removing the floating objects in the mesh, as shown in Fig.
+With the help of high-quality renderings, the object mask is extracted, and the depth image is masked.
+The noise in the depth image can be filtered out, thereby removing the floaters in the mesh, as shown in Fig.
 7 (b).
+Experiments show that the proposed depth filtering method is beneficial for removing floaters and obtaining a cleaner mesh.
 Loss function.
@@
 As shown in Fig.
-7, the two regularization terms in the loss function are important for extracting high-quality meshes.
-The Normal Consistency regularization term can make the normals of Gaussians more accurate, and the Depth Distortion term can make the Gaussians distributed on the plane.
+8, the two regularization terms in the loss function are important for extracting high-quality meshes.
+Among them, the normal consistency regularization term can make the normal of the Gaussian sphere more accurate, thereby making the surface of the mesh smoother.
+The Depth Distortion term can make the Gaussians distributed on the plane, thus reducing the surface damage.
 The corresponding metrics are shown in Tab.
-4.
-Limitations and Future work
+5.
+It can be found from the table that adding two regularization terms to the loss function significantly reduces the CD value and improves the quality of the mesh.
+Figure 8. Ablation Study on loss function.
+Post-processing.
+Since the input view of the model is sparse, the extracted mesh is prone to holes in some areas that are difficult to see.
+Therefore, we use some post-processing methods to repair the holes that may appear in the mesh to improve the integrity of the mesh.
+As shown in Fig.
+9, we used MeshLab to repair the holes in the mesh.
+As can be seen from the figure, simple post-processing operations can fill in the holes that may appear in the reconstructed mesh.
+Figure 9. Repair holes in the mesh.
+$downarrow Method CD\downarrow EMD\downarrow$
+w/o DF 0.425 0.110 w/o DD 0.350 0.109
+w/o NC 0.443 0.101 Full 0.328 0.110
+Table 5. Ablation experiment results of the “bird” sample in the DGMesh dataset. (DF: Depth filtering, NC: Normal Consistency, $D$: Depth Distortion)
+6. Limitations and Future work
 Though dynamic 2D Gaussians can model the accurate dynamic surface, there are several limitations: (1) Some holes exist in the extracted mesh.
-(2) Given the extremely sparse views, such as strict monocular input, robustness priors can other low-rank motion representations (Wang et al.
-2024a) could be explored to model more accurate motion.
-(3) More post-processing methods could be extended to handle Geometry floaters and holes completion.
-Figure 7: Ablation Study on loss function.
-Conclusion
+We need some post-processing operations to repair these holes.
+Future work will pursue direct hole-free generation to bypass repair entirely.
+(2) Under ultra-sparse (e.g., monocular) views, robust priors and low-rank motion can enhance accuracy.
+Recently, 3D foundation models (Xiang et al., 2025; Tochilkin et al., 2024; Li et al., 2025; Ye et al., 2025) show the great potential to generate high quality mesh from a static image.
+We believe that reconstruct high quality mesh from the dynamic multiview input will be coming soon.
+7. Conclusion
 In this paper, we propose Dynamic 2D Gaussians to extract high-quality meshes from dynamic 2D images input.
 We use 2D Gaussians as the basic representation primitives and use sparse control points to control 2D Gaussians to model the motion of objects.
-We used high-quality rendered images to filter out the noise in the depth image and used TSDF to obtain higher-quality meshes.
+We extract masks from high-quality rendering data and use these masks to filter out noise in the depth map.
+Finally, we use TSDF to extract the mesh of the object at any time.
 Experiments show that the mesh extracted by the proposed method has high details and a smoother surface.
-Appendix
-Written by AAAI Press Staff1With help from the AAAI Publications Committee.
-AAAI Style Contributions by Pater Patel Schneider, Sunil Issar,
-[…]
-Figure 8: Relighting and composition results in generated mesh on the Blender.
-Figure 9: Rendered images, normals, and depth maps.
-Figure 10: Comparison of results with different background colors.
-In the supplementary material, we mainly provide additional experimental results in Sec.
-1.
-Then more discussions are conducted in Sec.
-2.
-Finally, we delve into the limitations of our proposed Dynamic 2D Gaussians in Sec.
-3.
-Additional Experimental Results
-In the D-NeRF dataset, the metrics of the RGB images rendered from the extracted meshes do not fully reflect the quality of the meshes.
-Therefore, we provide additional visualization results, as shown in Fig.
-11.
-DG-Mesh exhibits relatively poorer performance in modeling object details.
-The meshes extracted by Deformable-3DGS and SCGS exhibit numerous pits and holes on the surface.
-In comparison, the meshes extracted by the proposed Dynamic 2D Gaussians approach capture object details while maintaining a smoother surface, reflecting better geometric properties.
-Figure 11: Mesh and mesh rendering visualization results of the D-NeRF dataset.
-In the DG-Mesh dataset, the detailed metrics are shown in Table 1.
-Since the data of the DGMesh dataset is relatively simple and the surface is smooth, it is difficult to reflect the advantages of the proposed method in terms of metrics.
-Our metrics are close to those of SCGS, but the visual quality shows that our extracted surfaces appear smoother than SCGS.
-The DG-Mesh method shows advantages in the Duck and Torus2sphere examples, but performs relatively poorly on the other cases.
-Although DGMesh has a high metrics in Torus2sphere, from the visualization results in Fig.
-12, DGMesh does not show a particularly good reconstruction effect in the Torus2sphere example.
-The proposed method can render Gaussians to obtain normal depth maps and high-quality RGB images of dynamic objects, as shown in Fig. 9.
-Method Duck Horse Bird
-[…]
-D-3DGS 1.064 0.076 0.298 0.114 0.427 0.123
-SCGS 1.001 0.074 0.232 0.118 0.336 0.121
-DG-mesh 0.782 0.048 0.299 0.168 0.557 0.128
-Ours 1.040 0.092 0.391 0.177 0.328 0.110
-Method Beagle Torus2sphere Girlwalk
-[…]
-D-3DGS 0.585 0.101 2.655 0.155 0.340 0.122
-SCGS 0.528 0.100 2.716 0.155 0.321 0.124
-DG-mesh 0.626 0.114 1.515 0.126 0.406 0.153
-Ours 0.544 0.122 2.479 0.164 0.324 0.129
-Table 5: Mesh reconstruction quality comparison results of DG-Mesh dataset.
-Figure 12: Visualization results of DGMesh dataset.
-Discussion and Application
-Background Color.
-Training models with different background colors can have some influence on the extraction of the mesh.
-The background color should be selected to have significant contrast with the color of the object itself.
-For the training images, the background color should be selected to have significant contrast with the color of the object itself.
-As shown in Fig.
-10, if a black background is used, the shadows on the object are difficult to distinguish from the background, making it challenging to extract the object’s mask from the rendered images.
-This then makes it difficult to filter out floating artifacts in the depth maps, ultimately leading to the presence of many floating artifacts in the extracted meshes.
-Mesh composition and relighting.
-The mesh reconstructed by Dynamic 2D Gaussians can be applied in the Blender in the Fig.
-8.
-Placing a proper light source, the shadow exists in the image due to the occlusion relationship between the objects.
-Time-Aware Mesh.
-Our method supports the extraction of dynamic mesh sequences, as shown in Fig.
-13.
-The extracted precise temporal geometry provides the potential for analyzing the motion characteristics and dynamics of objects.
-Figure 13: Time-Aware Mesh.
-Limitations
-Due to the thin nature of 2DGS and the sparsity of dynamic scene views, the extracted mesh may be broken in some areas.
-As shown in Fig.
-14, the mesh is prone to damage in the following two situations: 1.
-In some places that are difficult to observe from the training perspective and in shadows, the mesh is prone to damage, such as Fig.
-14 (a).
-2.
-Mesh is prone to damage on uneven thin planes with uneven lighting, such as Fig.
-14 (b).
-Subsequent research is needed on better methods for representing precise dynamic geometry, as well as targeted mesh damage repair techniques.
-This will help us better extract mesh from dynamic objects.
-Figure 14: Some areas of the mesh will appear broken.
+Acknowledgements.
+This work was supported by National Natural Science Foundation of China (No. 62376102).